OpenStudy (anonymous):

Solve 2x^2 = 5x + 4 using the quadratic formula. Round your answer to the nearest tenth. (Points : 1) {3.1, -0.6} {-3.1, 0.6} {1.9, -0.6} {-1.9, 0.6}

OpenStudy (nathan917):

Simplifying 2x2 = 5x + 4 Reorder the terms: 2x2 = 4 + 5x Solving 2x2 = 4 + 5x Solving for variable 'x'. Reorder the terms: -4 + -5x + 2x2 = 4 + 5x + -4 + -5x Reorder the terms: -4 + -5x + 2x2 = 4 + -4 + 5x + -5x Combine like terms: 4 + -4 = 0 -4 + -5x + 2x2 = 0 + 5x + -5x -4 + -5x + 2x2 = 5x + -5x Combine like terms: 5x + -5x = 0 -4 + -5x + 2x2 = 0 Begin completing the square. Divide all terms by 2 the coefficient of the squared term: Divide each side by '2'. -2 + -2.5x + x2 = 0 Move the constant term to the right: Add '2' to each side of the equation. -2 + -2.5x + 2 + x2 = 0 + 2 Reorder the terms: -2 + 2 + -2.5x + x2 = 0 + 2 Combine like terms: -2 + 2 = 0 0 + -2.5x + x2 = 0 + 2 -2.5x + x2 = 0 + 2 Combine like terms: 0 + 2 = 2 -2.5x + x2 = 2 The x term is -2.5x. Take half its coefficient (-1.25). Square it (1.5625) and add it to both sides. Add '1.5625' to each side of the equation. -2.5x + 1.5625 + x2 = 2 + 1.5625 Reorder the terms: 1.5625 + -2.5x + x2 = 2 + 1.5625 Combine like terms: 2 + 1.5625 = 3.5625 1.5625 + -2.5x + x2 = 3.5625 Factor a perfect square on the left side: (x + -1.25)(x + -1.25) = 3.5625 Calculate the square root of the right side: 1.887458609 Break this problem into two subproblems by setting (x + -1.25) equal to 1.887458609 and -1.887458609. Subproblem 1 x + -1.25 = 1.887458609 Simplifying x + -1.25 = 1.887458609 Reorder the terms: -1.25 + x = 1.887458609 Solving -1.25 + x = 1.887458609 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '1.25' to each side of the equation. -1.25 + 1.25 + x = 1.887458609 + 1.25 Combine like terms: -1.25 + 1.25 = 0.00 0.00 + x = 1.887458609 + 1.25 x = 1.887458609 + 1.25 Combine like terms: 1.887458609 + 1.25 = 3.137458609 x = 3.137458609 Simplifying x = 3.137458609 Subproblem 2 x + -1.25 = -1.887458609 Simplifying x + -1.25 = -1.887458609 Reorder the terms: -1.25 + x = -1.887458609 Solving -1.25 + x = -1.887458609 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '1.25' to each side of the equation. -1.25 + 1.25 + x = -1.887458609 + 1.25 Combine like terms: -1.25 + 1.25 = 0.00 0.00 + x = -1.887458609 + 1.25 x = -1.887458609 + 1.25 Combine like terms: -1.887458609 + 1.25 = -0.637458609 x = -0.637458609 Simplifying x = -0.637458609 Solution The solution to the problem is based on the solutions from the subproblems. x = {3.137458609, -0.637458609}

OpenStudy (nathan917):

Id say A.